The length of a conference room is one and a half times its width.A carpet that is twice as long as it is wide is placed in the center of the room, leaving a 3m wide border round the carpet.Find the area of the carpet.
Answer by edjones(8007)

Let the width of the room, w, = x and the length, L, = 1.5x
The width of the carpet is, then, x-6 and the length of the carpet is 2(x-6).
Length of carpet + 6 = length of room: 2(x-6)+6=1.5x
2x-12+6=1.5x
.5x=6
x=12 m length of room
18 m width of room
12-6 by 18-6
=12 by 6 dimensions of carpet
12*6=72 m^2 Area of carpet.
.
Ed